OpenStudy (anonymous):

10.15] Sharmi and Reese each deposit money into a savings account. Sharmi deposits $200 less than Reese. After the same amount of time, Reese has earned $90 in interest at a rate of 3% per year, and Sharmi has earned $100 in interest at a rate of 5% per year. How much money did Reese and Sharmi initially deposit into their savings accounts?

OpenStudy (anonymous):

Reese initially deposited $400, Sharmi initially deposited $200. Reese initially deposited $600, Sharmi initially deposited $200. Reese initially deposited $400, Sharmi initially deposited $600. Reese initially deposited $600, Sharmi initially d

OpenStudy (anonymous):

Sorry, revised answers. :) Reese initially deposited $400, Sharmi initially deposited $200. Reese initially deposited $600, Sharmi initially deposited $200. Reese initially deposited $400, Sharmi initially deposited $600. Reese initially deposited $600, Sharmi initially deposited $400

OpenStudy (amistre64):

its always a good idea to rewrite the question to weed out the useless parts Sharmi deposits $200 less than Reese. S = R-200 After the same amount of time, Reese has earned $90 in interest at a rate of 3% per year, R(.03) = 90 Sharmi has earned $100 in interest at a rate of 5% per year S(.05) = 100 How much money did Reese and Sharmi initially deposit into their savings accounts?

OpenStudy (amistre64):

since we see that S = R-200, sub that into the S(.05) = 100 and solve for R

OpenStudy (amistre64):

if anything, the interest part can get more complicated; depending on if its simple interest or not

OpenStudy (anonymous):

Yeah, I guess. :(

OpenStudy (anonymous):

Uh..174.8?

OpenStudy (amistre64):

I just read up on the simple interest stuff, and i believe my setup is correct

OpenStudy (amistre64):

S(.05) = 100 S = R-200 (R-200)(.05) = 100 R(.05) - 200(.05) = 100 R(.05) = 100+ 200(.05) R = (100+ 200(.05))/(.05)

OpenStudy (anonymous):

2200?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Oh,ok, thx, but that isn't a multiple choice answer?

OpenStudy (amistre64):

yeah, you sure you got the values correct?

OpenStudy (amistre64):

S = 100/.05 = 2000

OpenStudy (anonymous):

oh, okay. :) but tht is not a multiple choice either? the only ones are these... Reese initially deposited $400, Sharmi initially deposited $200. Reese initially deposited $600, Sharmi initially deposited $200. Reese initially deposited $400, Sharmi initially deposited $600. Reese initially deposited $600, Sharmi initially deposited $400

OpenStudy (anonymous):

Reese initially deposited $400, Sharmi initially deposited $600.

OpenStudy (anonymous):

Oh, okay, thx, do you have any work?

OpenStudy (amistre64):

becasue 600 is 200 less than 400??

OpenStudy (anonymous):

That you could show me?

OpenStudy (anonymous):

Oh, okay, got it! I see if I hget it right. :(

OpenStudy (amistre64):

its either going to be A or D in this case, since those are the only 2 options were R is greater than S

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

it is C

OpenStudy (anonymous):

i am sure

OpenStudy (anonymous):

Sharmi deposits $200 less than Reese

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

It might be c, but also a, I am so confused!

OpenStudy (amistre64):

C is Reese initially deposited $400, Sharmi initially deposited $600. R is less than S , which is simply not applicable to the problem that was stated

OpenStudy (anonymous):

so is c is right?

OpenStudy (amistre64):

is 600 less than 400 ????

OpenStudy (anonymous):

Well, duh? But 200 is less than 400 too!!!

OpenStudy (amistre64):

Reese minus 200 equals Sharmi so its either A or D, since those are the only options where Reese has more than Sharmi

OpenStudy (anonymous):

yeah, I guess! Which one????

OpenStudy (amistre64):

how many chances you got?

OpenStudy (anonymous):

After this test, it is like my last chance!

OpenStudy (anonymous):

Do you know anyone who knows Algebra?

OpenStudy (amistre64):

yeah, me

OpenStudy (amistre64):

and the options dont match the algebra

OpenStudy (anonymous):

ohh...okay, which one do you think? this is Algebra 1 !!!

OpenStudy (amistre64):

which one do they want you to choose? I dunno, none of them match the question that is asked. But If i were to take a guess, it would be a toss up between A and D

OpenStudy (anonymous):

I closed this question a little early so, yeah...:)

OpenStudy (anonymous):

I guess I wll try D, that's what katragaddasaichandra said. :)

OpenStudy (anonymous):

Sorry, C. :)

OpenStudy (amistre64):

Katra said C Sharmi has earned $100 in interest at a rate of 5% n(.05) = 100; when n = 2000 and you dont have any options that have 2000 as a solution

OpenStudy (amistre64):

can you take a screen shot by chance so that we could see that actual problem?

OpenStudy (anonymous):

That's the exact way the ? is written on my test !!!!

OpenStudy (anonymous):

Should I just go with C?

OpenStudy (amistre64):

I spose you could go with C, it makes as much sense as the other options ... the only thing is that it doesnt conform to the rule that Reese has 200 MORE THAN Sharmi.

OpenStudy (anonymous):

Yeah, well thx, for your help! I will let you know if I get it right or wrong. :(

OpenStudy (amistre64):

if you copy and paste the first sentence of the problem into google, you get similar if not the same problem; such as this one http://openstudy.com/study#/updates/4fedfd03e4b0bbec5cfd410c

OpenStudy (anonymous):

Ok, I will

OpenStudy (anonymous):

Hmm...not very helpful!!!

OpenStudy (amistre64):

i think i see something that i was missing

OpenStudy (amistre64):

i assumed the time to be a year; and it is not specified

OpenStudy (anonymous):

Oh..okay. :D

OpenStudy (amistre64):

\[S=R-200\]\[100=S(.05)(t)\]\[90=R(.03)(t)\] \[\frac{100}{.05t}=S\]\[\frac{90}{.03t}=R\] \[\frac{100}{.05t}=\frac{90}{.03t}-200\] \[\frac{100}{.05}=\frac{90}{.03}-200t\] \[\large \frac{\frac{100}{.05}-\frac{90}{.03}}{-200}=t\]

OpenStudy (anonymous):

So the answer is c?

OpenStudy (amistre64):

t = 5 ... which .... no, the answer is not C since 400 is NOT less than 600, still

OpenStudy (anonymous):

nope, you r right! W=either a or d. :D Do you know which one?

OpenStudy (amistre64):

now that i know t, i can solve for S or R\[\frac{100}{.05t}=S\]\[\frac{100}{5(.05)}=S\]\[\frac{20}{.05}=S\] so what is S according to this?

OpenStudy (anonymous):

400/

OpenStudy (amistre64):

then Sarmi has 400 and Reese has 200 more, or 600; does D match that?

OpenStudy (anonymous):

Yes1 OMG! Thank you! I am really bad at Algebra! LOL> :)

OpenStudy (amistre64):

'sok, that "time" was the key. I had stupidly assumed it to be 1

OpenStudy (anonymous):

LOL! Can you help me with one more?

OpenStudy (amistre64):

write it up in a new post, and tag me in it so i can find it

OpenStudy (amistre64):

i might be able to work something out

OpenStudy (anonymous):

Ok, thx, the new post is coming upo, try to find it! :)

OpenStudy (amistre64):

im old and blind .... use: @amistre64 to shot me a link to it